OpenStudy (anonymous):

Simplify the following: (x^2 - 1)^2 * Sqrt(x + 1) -------------------- (x - 1)^3/2 Can anyone show me the steps to get to the result: (x + 1)^2 * Sqrt(x^2 - 1) I can't figure it out.

OpenStudy (anonymous):

\[\frac{\left(x^2-1\right)^{2 }(x+1)^{1/2}}{(x-1)^{\frac{3}{2}}}\] Multiply Top and Bottom by \[(x-1)^{-\frac{3}{2}}\]

OpenStudy (anonymous):

\[(x^2 - 1)^2 ==(x+1)^2(x-1)^2\] \[(x-1)^{-3/2}\left((x+1)^2(x-1)^2(x+1)^{1/2}\right)\]

OpenStudy (anonymous):

\[(x+1)^{5/2}(x-1)^2(x-1)^{-3/2}\] \[(x-1)^{1/2}(x+1)^{5/2}\]